# HistDiST: Histopathological Diffusion-based Stain Transfer

### v1

@@ Abstract
Hematoxylin and Eosin (H&E) staining is the cornerstone of histopathology but lacks molecular specificity. While Immunohistochemistry (IHC) provides molecular insights, it is costly and complex, motivating H&E-to-IHC translation as a cost-effective alternative. Existing translation methods are mainly GAN-based, often struggling with training instability and limited structural fidelity, while diffusion-based approaches remain underexplored. We propose HistDiST, a Latent Diffusion Model (LDM) based framework for high-fidelity H&E-to-IHC translation. HistDiST introduces a dual-conditioning strategy, utilizing Phikon-extracted morphological embeddings alongside VAE-encoded H&E representations to ensure pathology-relevant context and structural consistency. To overcome brightness biases, we incorporate a rescaled noise schedule, $v$-prediction, and trailing timesteps, enforcing a zero-SNR condition at the final timestep. During inference, DDIM inversion preserves the morphological structure, while an $\eta$-cosine noise schedule introduces controlled stochasticity, balancing structural consistency and molecular fidelity. Moreover, we propose Molecular Retrieval Accuracy (MRA), a novel pathology-aware metric leveraging GigaPath embeddings to assess molecular relevance. Extensive evaluations on MIST and BCI datasets demonstrate that HistDiST significantly outperforms existing methods, achieving a 28% improvement in MRA on the H&E-to-Ki67 translation task, highlighting its effectiveness in capturing true IHC semantics.
Keywords:
Stain Transfer Diffusion Models Digital Pathology
@@ 1 Introduction
Histopathology is essential for disease diagnosis, revealing tissue structure and cellular morphology. While Hematoxylin and Eosin (H&E) staining is cost-effective and highlights structural details, it lacks molecular specificity. In contrast, Immunohistochemistry (IHC) provides molecular insights but is expensive and time-consuming. Automated H&E-to-IHC translation offers a scalable solution by inferring molecular details from tissue morphology, reducing costs, and enhancing diagnostic consistency. GANs have shown promise in this task. Li et al. [2] use adaptive supervised PatchNCE loss to preserve structure, while MDCL [10] employs a conditional GAN with multi-domain contrastive learning to enhance cross-domain alignment. Pathology-specific approaches include patch-level feature extraction with multiple instance learning [4] and HER2 scoring enhancement via nuclei density estimation [9]. However, GANs still face mode collapse and require complex loss designs to maintain consistency in pathology screenings.
Recently, diffusion models (DMs) [18] have emerged as a powerful alternative, offering superior stability and synthesis quality. Latent Diffusion Models (LDMs) [19] reduce the computational burden of DMs by operating in the latent space of pretrained autoencoders. LDM-based image editing methods, such as InstructPix2Pix [21] for instruction-based editing and ControlNet [8] for conditional generation, offer structured control. Inference-time optimization techniques using DDIM Inversion [23] and Delta Denoising Score [24] further enhance structural fidelity. Although there has been quite some work on applying LDMs for natural image editing, the application of LDMs for H&E-to-IHC stain translation remains largely unexplored. Existing diffusion-based methods [12][13] mainly use text prompts or unpaired signal vector conditioning, but often fail to ensure structural and semantic consistency, critical for accurate pathology translation.
@@
HistDiST overcomes these challenges through three key contributions. First, a dual-conditioning strategy integrates Phikon-based [20] pathology-specific embeddings into the U-Net’s cross-attention layers and concatenates VAE-encoded H&E representations with the input latent noise. This ensures that pathology-relevant context guides denoising for precise molecular reconstruction. Second, to address brightness inconsistencies, we incorporate a rescaled noise schedule and trailing timesteps [26] along with $v$-prediction [25], enforcing zero terminal SNR. Third, we perform joint training for unconditional H&E generation and H&E-to-IHC translation, a critical step that enables the use of DDIM inversion [16] for structure preservation at inference. However, the deterministic nature of DDIM sampling restricts generative flexibility, limiting the model’s ability to capture subtle molecular variations critical for accurate IHC synthesis. To mitigate this, we incorporate an $\eta$-cosine noise schedule, progressively increasing the noise scale during denoising to enable stochastic exploration. This significantly improves FID scores while preserving both morphological structures and molecular fidelity in the translated IHC images. Moreover, we propose a novel feature-based metric, Molecular Retrieval Accuracy (MRA), that assesses molecular fidelity via cosine similarity. Experiments on the MIST and BCI datasets demonstrate that HistDiST outperforms state-of-the-art methods across both qualitative and quantitative metrics, achieving superior visual fidelity, structural preservation, and molecular fidelity. Notably, HistDiST shows 28% improvement on MRA on H&E-to-Ki67, demonstrating highly accurate IHC synthesis.
@@ 2.1 Preliminaries
Diffusion models generate data by learning to reverse a stochastic forward process that gradually corrupts an input sample $x_{0}$ with Gaussian noise over $T$ steps, following a variance schedule $\beta_{t}$: $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$. For large $T$, this process converges to an isotropic Gaussian distribution. The model approximates the reverse process, $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ where $\mu_{\theta}$ and $\Sigma_{\theta}$ are learnable functions. Training involves predicting the noise component $\epsilon_{\theta}(x_{t},t)$ by minimizing: $L=\mathbb{E}_{x_{0},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$, where $\epsilon\sim\mathcal{N}(0,I)$ and $\epsilon_{\theta}$ is a neural network parameterized by $\theta$. Latent Diffusion Models (LDMs) enhance computational efficiency by operating in a compressed latent space, where an encoder $E$ maps images $x_{0}$ to latent representations $z_{0}=E(x_{0})$.
@@ 2.2 HistDiST
We present HistDiST, a novel framework leveraging LDM for high-fidelity H&E-to-IHC stain translation. LDM operate in a variational autoencoder (VAE) latent space and use a U-Net denoising network with residual connections, self-attention, and cross-attention mechanisms. HistDiST achieves superior structural and molecular accuracy through: (1) dual-conditioning (2) v-prediction with rescaled noise schedules, and (3) DDIM inversion with $\eta$-cosine scheduling balancing determinism with controlled stochasticity.
@@
Moreover, when SNR is zero, traditional epsilon ($\epsilon$) prediction becomes trivial and does not provide meaningful learning signals. To overcome this, HistDiST adopts a $v$-prediction and $v$-loss framework as proposed in [25]. The velocity vector is defined as $v_{t}=\sqrt{\bar{\alpha}_{t}}\,\epsilon-\sqrt{1-\bar{\alpha}_{t}}\,x_{0}$, where $x_{0}$ is the clean latent representation (noise-free image), $\epsilon$ is the Gaussian noise added during the forward diffusion process, and $\bar{\alpha}_{t}$ represents the cumulative product of the noise schedule parameters up to timestep $t$. The corresponding v-prediction loss function is:
$\mathcal{L}=\lambda_{t}\|v_{t}-\tilde{v}_{t}\|_{2}^{2}$ (1)
@@
$z_{t+1}=\sqrt{\alpha_{t+1}}\left(\frac{z_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}(z_{t},t,c_{T})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t+1}}\,\epsilon_{\theta}(z_{t},t,c_{T})$ (2)
where $\alpha_{t}$ is the cumulative noise schedule, $\epsilon_{\theta}$ is the predicted noise, and $c_{T}$ is the text-based conditioning embedding. While using this inverted noise as the starting point for translation successfully preserved structural features, we observed a increase in FID scores (Table 2 (right), $\eta$=0), indicating that deterministic DDIM sampling restricted generative flexibility and hindered the model’s ability to capture molecular diversity essential for accurate IHC synthesis.
@@
$\eta_{\text{t}}(t)=0.6-0.4\cos\left(\pi\frac{t}{T}\right)$
@@
$z_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{z_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(z_{t},t)}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}(z_{t},t)+\sigma_{t}\cdot\eta_{t}\cdot\epsilon$ (3)
@@ 3 Experiments
$\mathrm{MIST}_{ER}$ CycleGAN 0.1982 0.5175 0.5092 0.3710 0.8672 0.235 125.7 95.1
@@
HistDiST 0.2278 0.3645 0.3105 0.2465 0.8226 0.552 31.2 3.6
$\mathrm{MIST}_{HER2}$ ASP 0.2004 0.4534 0.4150 0.2665 0.8174 0.193 51.4 12.4
@@ 3.0.2 Implementation Details:
We fine-tune Stable Diffusion v1.5, training for 500 epochs with a batch size of 16. AdamW optimization is used with a $42e-$ 4 learning rate, 1000 warmup steps, and a cosine decay schedule. Training alternates between H&E generation and H&E-to-IHC translation. $v$-prediction is set to $\gamma=5$, and noise schedule is rescaled to enforce zero terminal SNR.
@@ 3.0.4 Results:
We compare HistDiST with state-of-the-art stain translation models—CycleGAN [5], Pix2Pix [6], PyramidP2P [7], ASP [2], and MDCL [10]—for H&E-to-ER translation. For other transfer tasks, we select the top two performers for comparison. As shown in Table 1, HistDiST achieves consistent performance gains across all key metrics. It achieves the highest MRA scores, surpassing MDCL by 23% on MISTER and MISTHER2, demonstrating superior molecular fidelity. It also attains the lowest PHV scores, indicating better feature-space alignment, and the best FID/KID scores, confirming that the generated IHC images are more visually realistic and distributionally aligned with real ones.
3.0.5 Ablation Studies:
Table 2: Ablation study (on ER) during training (left) and inference (right). Abbreviations: "SD+IC": Stable Diffusion with input conditioning based on VAE-extracted H&E features, "$v$-pred": $v$-prediction, "Inv": DDIM Inversion with $\eta$-cosine start $20\%$ noise (for DDIM sampling) , "xAtt": cross-attention with Phikon-extracted H&E features. The inference strategies involve DDIM inversion for structure preservation, followed by DDIM sampling with different noise schedules (illustrated in Fig. 2 (right))

### v2

@@ Abstract
Hematoxylin and Eosin (H&E) staining is the cornerstone of histopathology but lacks molecular specificity. While Immunohistochemistry (IHC) provides molecular insights, it is costly and complex, motivating H&E-to-IHC translation as a cost-effective alternative. Existing translation methods are mainly GAN-based, often struggling with training instability and limited structural fidelity, while diffusion-based approaches remain underexplored. We propose HistDiST, a Latent Diffusion Model (LDM) based framework for high-fidelity H&E-to-IHC translation. HistDiST introduces a dual-conditioning strategy, utilizing Phikon-extracted morphological embeddings alongside VAE-encoded H&E representations to ensure pathology-relevant context and structural consistency. To overcome brightness biases, we incorporate a rescaled noise schedule along with $v$-prediction, enforcing a zero-SNR condition at the final timestep. During inference, DDIM inversion preserves the morphological structure, while an $\eta$-cosine noise schedule introduces controlled stochasticity, balancing structural consistency and molecular fidelity. Moreover, we propose Molecular Retrieval Accuracy (MRA), a novel pathology-aware metric leveraging GigaPath embeddings to assess molecular relevance. Extensive evaluations on MIST and BCI datasets demonstrate that HistDiST significantly outperforms existing methods, achieving a 28% improvement in MRA on the H&E-to-Ki67 translation task, highlighting its effectiveness in capturing true IHC semantics. The code is available at https://github.com/ErikGro/HistDiST.
@@ 1 Introduction
Histopathology is essential for disease diagnosis, revealing tissue structure and cellular morphology. While Hematoxylin and Eosin (H&E) staining is cost-effective and highlights structural details, it lacks molecular specificity. In contrast, Immunohistochemistry (IHC) provides molecular insights but is expensive and time-consuming. Automated H&E-to-IHC translation offers a scalable solution by inferring molecular details from tissue morphology, reducing costs, and enhancing diagnostic consistency. Deep learning methods, particularly GANs, GANs have shown promise in this task. Li et al.  [11] use adaptive supervised PatchNCE loss to preserve structure, while MDCL  [23] employs a conditional GAN with multi-domain contrastive learning to enhance cross-domain alignment. Pathology-specific approaches include patch-level feature extraction with multiple instance learning  [12] and HER2 scoring enhancement via nuclei density estimation  [17]. However, GANs still face mode collapse and require complex loss designs to maintain consistency in pathology screenings.
Recently, diffusion models (DMs)  [8] have emerged as a powerful alternative, offering superior stability and synthesis quality. Latent Diffusion Models (LDMs)  [19] reduce the computational burden of DMs by operating in the latent space of pretrained autoencoders. LDM-based image editing methods, such as InstructPix2Pix  [1] for instruction-based editing and ControlNet  [26] for conditional generation, offer structured control. Inference-time optimization techniques using DDIM Inversion [22] and Delta Denoising Score [7] further enhance structural fidelity. Although there has been quite some work on applying LDMs for natural image editing, the application of LDMs for H&E-to-IHC stain translation remains largely unexplored. Existing diffusion-based methods [4, 25] mainly use text prompts or unpaired signal vector conditioning, but often fail to ensure structural and semantic consistency, critical for accurate pathology translation.
@@
HistDiST overcomes these challenges through three key contributions. First, we employ a dual-conditioning strategy that integrates pathology-specific priors into denoising process. Specifically, Phikon-based embeddings [5] are injected via cross-attention within the U-Net, while VAE-encoded H&E representations are concatenated with latent noise input. This ensures that pathology-relevant context guides denoising for precise molecular reconstruction. Second, to address brightness inconsistencies, we incorporate a rescaled noise schedule and trailing timesteps  [13] along with $v$-prediction  [10], enforcing zero terminal SNR. Third, we perform joint training for unconditional H&E generation and H&E-to-IHC translation, a critical step that enables the use of DDIM inversion  [21] for structure preservation at inference. However, while inversion effectively anchors structure, the subsequent DDIM denoising—being fully deterministic—restricts generative flexibility. As a result, the model fails to capture subtle but diagnostically relevant molecular variations in the translated IHC image.
To overcome this limitation, we introduce an $\eta$-cosine noise schedule that progressively increases stochasticity during DDIM denoising. This modification enables controlled exploration of molecular solution space while preserving morphological consistency. The resulting outputs show improvements in synthesis diversity and fidelity, as reflected by lower FID scores. Moreover, we propose Molecular Retrieval Accuracy (MRA), a novel feature-based metric that quantifies molecular fidelity using cosine similarity. Experiments on MIST and BCI datasets demonstrate that HistDiST outperforms state-of-the-art methods across both qualitative and quantitative benchmarks, achieving superior visual realism, structural preservation, and molecular accuracy. Notably, HistDiST yields 28% improvement in MRA on H&E-to-Ki67 translation, demonstrating highly accurate IHC synthesis.
2 Related Work
2.1 Image Generation using Diffusion Models
Diffusion models (DMs) have emerged as a powerful alternative to GANs for image generation, offering greater diversity and training stability by learning the data distribution through iterative denoising processes [8]. These models operate either directly in the data space or in a compressed latent space, as in Latent Diffusion Models (LDMs)[19], enabling more efficient computation. Building on their success in generative tasks, conditional diffusion models have been developed to allow guided image synthesis for applications such as image-to-image translation [16] and text-based image editing  [1]. Conditioning is typically achieved by injecting external information into the denoising network. For instance, Stable Diffusion [19] and InstructPix2Pix [1] leverage text prompts mapped to a shared embedding space with images, while ControlNet [27] enhances this framework by incorporating spatial guidance, enabling fine-grained structural control. In addition, inference-time optimization techniques—such as DDIM inversion [22] and Delta Denoising Score (DDS) [7]—have been introduced to improve structural fidelity and control in a training-free manner.
2.2 Stain Transfer in Computational Pathology
The generation of virtually stained images given a source staining type image is an active field of research, aiming to alleviate the lack of useful IHC counterparts to H&E images and meanwhile preventing the challenges caused by re-staining or re-sampling tissues. Generative Adversarial Networks (GANs) have been widely employed for H&E-to-IHC translation, with many approaches focusing on preserving structural and pathological integrity. For example, Li et al.[11] introduce an adaptive supervised PatchNCE loss to maintain structure, while Chen et al.[3] improve translation fidelity via a multi-branch discriminator with prototype consistency. Other methods incorporate domain-specific enhancements: Li et al.[12] propose a patch-level pathological feature extractor with multiple-instance learning, and Peng et al.[17] refine HER2 scoring using a nuclei density estimator and auxiliary alignment branch. Wang et al.[23] incorporate Multi-Domain Contrastive Learning (MDCL) within a conditional GAN to enhance cross-domain alignment, and Qu et al.[18] utilize a multi-magnification processing strategy with an attention module to extract fine-grained features while minimizing information loss. Despite their effectiveness, GAN-based methods often struggle with mode collapse and require carefully crafted loss functions to ensure pathological accuracy.
Recently, diffusion models have gained attention for their stability and high-quality synthesis in natural image generation. Their application in virtual staining, however, remains limited. He et al.[6] leverage diffusion models augmented with Schrödinger bridges and structural constraints to achieve stain transfer, though standard evaluation metrics are not reported. Shen et al.[20] employ diffusion solely for style transfer, without addressing cross-stain learning. Text-conditioned diffusion has also been explored: Dubey et al.[4] utilize a latent diffusion model guided by text prompts, but their reliance on pixel-aligned datasets limits practical applicability. Yan et al.[25] propose a unified dual-encoder diffusion framework for multi-stain translation across diverse stain types (H&E, MT, PAS, PASM). Nonetheless, the use of diffusion models for H&E-to-IHC stain translation remains underexplored, with a lack of methods that explicitly target pathological consistency and structural fidelity in this challenging setting.
@@ 3.1 Preliminaries
Diffusion models generate data by learning to reverse a stochastic forward process that gradually corrupts an input sample $x_{0}$ with Gaussian noise over $T$ steps, following a variance schedule $\beta_{t}$:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I).$ (1)
For large $T$, this process converges to an isotropic Gaussian distribution. The model approximates the reverse process:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$ (2)
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are learnable functions. Training involves predicting the noise component $\epsilon_{\theta}(x_{t},t)$ by minimizing:
$L=\mathbb{E}_{x_{0},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right],$ (3)
where $\epsilon\sim\mathcal{N}(0,I)$ and $\epsilon_{\theta}$ is a neural network parameterized by $\theta$. Latent Diffusion Models (LDMs) enhance computational efficiency by operating in a compressed latent space, where an encoder $E$ maps images $x_{0}$ to latent representations $z_{0}=E(x_{0})$.
@@ 3.2 HistDiST
We present HistDiST, a novel framework leveraging LDMs for high-fidelity H&E-to-IHC translation. LDMs operate in a variational autoencoder (VAE) latent space and use a U-Net denoising network with residual connections, self-attention, and cross-attention mechanisms. HistDiST achieves superior structural and molecular accuracy through: (1) dual-conditioning (2) v-prediction with rescaled noise schedules, and (3) DDIM inversion followed by DDIM denoising with $\eta$-cosine scheduling, balancing determinism with controlled stochasticity.
@@
Moreover, when SNR is zero, traditional epsilon ($\epsilon$) prediction becomes trivial and does not provide meaningful learning signals. To overcome this, HistDiST adopts a $v$-prediction and $v$-loss framework as proposed in  [10]. The velocity vector is defined as:
$v_{t}=\sqrt{\bar{\alpha}_{t}}\,\epsilon-\sqrt{1-\bar{\alpha}_{t}}\,x_{0},$ (4)
where $x_{0}$ is the clean latent representation (noise-free image), $\epsilon$ is the Gaussian noise added during the forward diffusion process, and $\bar{\alpha}_{t}$ represents the cumulative product of the noise schedule parameters up to timestep $t$. The corresponding v-prediction loss function is:
$\mathcal{L}=\lambda_{t}\|v_{t}-\tilde{v}_{t}\|_{2}^{2},$ (5)
@@
$z_{t+1}=\sqrt{\alpha_{t+1}}\left(\frac{z_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}(z_{t},t,c_{T})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t+1}}\,\epsilon_{\theta}(z_{t},t,c_{T}),$ (6)
where $\alpha_{t}$ is the cumulative noise schedule, $\epsilon_{\theta}$ is the predicted noise, and $c_{T}$ is the text-based conditioning embedding. While using this inverted noise as the starting point for translation successfully preserved structural features, we observed a increase in FID scores (Table 2 (right), $\eta$=0), indicating that deterministic DDIM denoising restricted generative flexibility and hindered the model’s ability to capture molecular diversity essential for accurate IHC synthesis.
@@
$\eta_{\text{t}}(t)=0.6-0.4\cos\left(\pi\frac{t}{T}\right),$ (7)
@@
$z_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{z_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(z_{t},t)}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}(z_{t},t)+\sigma_{t}\cdot\eta_{t}\cdot\epsilon,$ (8)
@@ 4 Experiments
$\mathrm{MIST}_{ER}$
CycleGAN 0.1982 0.5175 0.5092 0.3710 0.8672 0.235 125.7 95.1
@@
InstructPix2Pix 0.1536 0.5122 0.4961 0.3601 0.8679 0.210 61.3 38.9
ControlNet 0.2112 0.4992 0.4479 0.3370 0.8522 0.130 38.0 9.1
HistDiST 0.2278 0.3645 0.3105 0.2465 0.8226 0.552 31.3 3.6
$\mathrm{MIST}_{HER2}$
ASP 0.2004 0.4534 0.4150 0.2665 0.8174 0.193 51.4 12.4
@@ 4.0.2 Implementation Details:
We fine-tune Stable Diffusion v1.5, training for 500 epochs with a batch size of 16. AdamW optimization is used with a 2e42e-4 learning rate, 1000 warmup steps, and a cosine decay schedule. Training alternates between H&E generation and H&E-to-IHC translation. $v$-prediction is set to $\gamma=5$, and noise schedule is rescaled to enforce zero terminal SNR. Lastly, during inference, we use 200 inversion steps and 200 denoising steps.
@@ 4.0.4 Results:
We compare HistDiST with state-of-the-art stain translation models—CycleGAN  [28], Pix2Pix  [9], PyramidP2P  [14], ASP  [11], and MDCL  [23]—for H&E-to-ER translation. As there are no specific diffusion-based approaches for H&E-to-IHC stain transfer, we adapt two image translation methods widely used in the natural domain, InstructPix2Pix  [1] and ControlNet  [27] and make comparisons with them. Specifically, we fine-tune InstructPix2Pix for our task, which is also used as the initialization for ControlNet. We train ControlNet with Canny edge map  [2] of the H&E input as the conditioning signal. For stain translation tasks on other staining types, we select the top two GAN-based performers, along with InstructPix2Pix and ControlNet, for comparison. As shown in Table 1, HistDiST achieves consistent performance gains across all key metrics. It achieves the highest MRA scores, surpassing MDCL by 23% on MISTER and MISTHER2, demonstrating superior molecular fidelity. It also attains the lowest PHV scores, indicating better feature-space alignment, and the best FID/KID scores, confirming that the generated IHC images are more visually realistic and distributionally aligned with real ones.
4.0.5 Ablation Studies and Qualitative Results:
Table 2: Ablation study (on ER) during training (left) and inference (right). Abbreviations: "SD+IC": Stable Diffusion with input conditioning based on VAE-extracted H&E features (InstructPix2Pix), "$v$-pred": $v$-prediction, "Inv": DDIM Inversion followed by DDIM sampling with $\eta$-cosine start $20\%$ noise, "xAtt": cross-attention with Phikon-extracted H&E features. The inference strategies involve DDIM inversion for structure preservation, followed by DDIM sampling with different noise schedules (Fig. 2 (right))